I need to find the distance between two points (8, -5) and (-1, -2) I keep coming out with 4.90 and its wrong.
Answer by stanbon(75887)

I need to find the distance between two points (8, -5) and (-1, -2)
-----
Formula:: d = sqrt[(x1-x2)^2 + (y1-y2)^2]
-------------------
d = sqrt[(8--1)^2 + (-5--2)^2]
-----
d = sqrt[9^2 + 3^2]
=====
d = sqrt(90)
d = 3sqrt(10) = 3*3.16 = 9.49
================
Cheers,
Stan H.